OpenStudy (jake124):

65 in =____ ft A. 6 5/12 B. 12 C. 5 5/12 D. 780 i think its B please help

OpenStudy (danjs):

for 12 inches = 1 foot \[\frac{ 65~~inch }{ 1 }*\frac{ 1~ft }{ 12~~inch }=\frac{ 65~~ ft}{ 12 }\]

OpenStudy (jake124):

so A THANK U SO MUCH !!!!!